A 4:1 molar mixture of he and ch4 kept in avessel at 20 bar  pressure .due to hole in vessel   gas mixture leaks out . whatb is composition of mixture effusing out intially

Grade:12th pass

1 Answers

Navkar jain
5 years ago
total mass =5
total pressure =20
partial pressure of He=4/5*20
=16
partial pressure of ch4=20 -16=4
molar mass of he=4                                  
molar mass of ch4=16
x[MCH4]/[MHE]1/2
ratio=8ratio 1